OpenStudy (anonymous):

A gasoline gauge initially read 1/3 full. when 10 gallons were added it read 3/4 full. How many more gallons are needed to fill the tank?

OpenStudy (anonymous):

by addding 10 gallons, you add 5/12 of a tank. 5/12= .42. you need to gain 1/4 of a tank in order to be full.

OpenStudy (tkhunny):

Please show some of your own work. You MUST have something to show us so we can help you more specifically.

OpenStudy (anonymous):

so for every 10 gallons, you gain .42 of a tank

OpenStudy (tkhunny):

1/3 (4/12) full takes x gallons 3/4 (9/12) full takes x+10 gallons 1/1 (12/12) full takes x + 10 + y gallons Has to be some sense in there, somewhere.

OpenStudy (anonymous):

and i agree with @tkhunny its hard to help with this problem without seeing some of your work

OpenStudy (mathmath333):

\(\large \color{black}{ \dfrac{t}{3}+10=\dfrac{3t}{4}\hspace{.33em}\\~\\ }\)

OpenStudy (mathmath333):

where t is the total capacity

OpenStudy (mathmath333):

\(\large \color{black}{ t=24\hspace{.33em}\\~\\ \dfrac{3t}{4}=18\hspace{.33em}\\~\\ }\) gallons more needed =\(\large \color{black}{ 24-18=6\quad \text{gallons}\hspace{.33em}\\~\\ }\)